Kids Web

A World Wide Web Digital Library for Schoolkids

THE WORLD WIDE WEB

The Web contains enormous amounts of information, and is growing incredibly rapidly. The ambitious goal of the World Wide Web project is to make available all of human knowledge through global computer networks acting as a vast distributed electronic library. However only a small fraction of that information is useful, intelligible and interesting to schoolchildren. Finding this information using standard Web digital libraries and search engines can be very time-consuming and frustrating, especially for schoolkids.

Ideally the students should be presented with a subset of the Web that is very simple to navigate, and contains information targeted at the K-12 level. This is the goal of Kids Web. Each subject section contains a list of links to information that is understandable and interesting to schoolkids. There are also links to external lists of material on each subject which more advanced students can browse for further information.

THE HISTORY AND FUTURE OF KIDS WEB

Kids Web grew out of a computer multimedia class based on the World Wide Web and Mosaic which was taught to 8th graders as part of the 1994 Young Scholars Program at Syracuse University. It is being developed as part of The Living Textbook project, which applies high-performance computing and communications to K-12 education, and is funded by New York State. By the end of 1994, a number of schools in the state will be connected to the Northeast Parallel Architectures Center (NPAC) at Syracuse University by NYNet, a fast optical fiber network using state-of-the-art ATM networking technology.

One of the problems with using the Web is that downloading information can be very slow, since currently the Information Superhighway more closely resembles a dirt road. One of our goals is to use Web caching technology to store the documents on Kids Web on a centralized server at NPAC. This will allow the information to be accessed much more rapidly by students in these schools.

In order for Kids Web to fulfill the goal of being a comprehensive digital library for kids, it must be managed in a distributed fashion, so that Web-enabled teachers (and students) from all over the world can suggest links to relevant Web documents, and each subject listing will be edited by educators who specialize in that field. If you are interested in being an editor for a subject section, or want to suggest a document to be linked to Kids Web, please email kidsweb@npac.syr.edu. Comments and inquiries are also welcome.

WHERE TO FIND KIDS WEB

A prototype of Kids Web is now available by directing your World Wide Web client (e.g. Mosaic or Netscape) to http://www.npac.syr.edu/textbook/kidsweb/.

Paul Coddington

Northeast Parallel Architectures Center Syracuse University kidsweb@npac.syr.edu